For a recessive X-linked trait there are 9 affected females in every 10,000 women. How many men in 10,000 are affected by the trait if the population is in Hardy-Weinberg equilibrium?
Answer #1

Answer:

The frequency of recissive X-linked trait = q^2 = 9/10000 = 0.0009

Frequency of recessive allele = q= SQRT of 0.0009 = 0.03

The X-linked recessive allele frequency = recessive X-linked phenotype frequency in male (as males have only one X-chromosome)

Therefore the frequency of recessive X-linked trait = 0.03 or 3%

3% out of 10000 = 300

Hence, 300 men may the have the trait out of 10000 people.